OpenStudy (anonymous):

Allie has a birthday present for her brother. The length of the box that the present comes in is two times longer than the width of the box. The height of the box is 5 inches and its volume is 160 cubic inches. What is the length and width of the box?

OpenStudy (anonymous):

LXBXH =160 L =2B & H = 5 LX1/2LX5=160 Solve for L U get : L = 8 and B = 4

OpenStudy (callisto):

let the width of the box = x the length of the box = 2x Volume = width x length x height 160 = 5 (x)(2x) 10x^2 =160 x^2 = 16 x =? Then you can find the length of the box :)